# Inductive Graph Representation Learning with Quantum Graph Neural Networks

### v1

@@ Abstract
Quantum Graph Neural Networks (QGNNs) present a promising approach for combining quantum computing with graph-structured data processing. While classical Graph Neural Networks (GNNs) are renowned for their scalability and robustness, existing QGNNs often lack flexibility due to graph-specific quantum circuit designs, limiting their applicability to a narrower range of graph-structured problems, falling short of real-world scenarios. To address these limitations, we propose a versatile QGNN framework inspired by the classical GraphSAGE approach, utilizing quantum models as aggregators. In this work, we integrate established techniques for inductive representation learning on graphs with parametrized quantum convolutional and pooling layers, effectively bridging classical and quantum paradigms. The convolutional layer is flexible, enabling tailored designs for specific problems. Benchmarked on a node regression task with the QM9 dataset, we demonstrate that our framework successfully models a non-trivial molecular dataset, achieving performance comparable to classical GNNs. In particular, we show that our quantum approach exhibits robust generalization across molecules with varying numbers of atoms without requiring circuit modifications, slightly outperforming classical GNNs. Furthermore, we numerically investigate the scalability of the QGNN framework. Specifically, we demonstrate the absence of barren plateaus in our architecture as the number of qubits increases, suggesting that the proposed quantum model can be extended to handle larger and more complex graph-based problems effectively.
1 Introduction
@@
Graphs provide a versatile framework for modeling complex relationships, with applications ranging from molecular structures to social networks. Unlike traditional machine learning models such as convolutional neural networks (CNNs) and feedforward neural networks (NNs), graph neural networks (GNNs) are specifically designed to exploit the relational structure inherent in graph data. This unique capability enables GNNs to excel in tasks such as node classification, link prediction, and graph regression, where understanding the relationships between entities is as critical as understanding the entities themselves. As a result, GNNs have found widespread use across diverse domains, including social network analysis [1, 2], recommendation systems [3, 4], drug discovery [5], traffic prediction [6], and polypharmacy side effects [7, 8].
@@
While classical GNNs have achieved remarkable success, the emergence of quantum computing introduces transformative potential for machine learning. Quantum Machine Learning (QML) harnesses principles of quantum mechanics—such as entanglement, superposition, and interference—to achieve significant, and in some cases exponential [9], speedups over classical computing for specific tasks [10, 11, 12, 13, 14]. In the context of graph-based problems, Quantum Graph Neural Networks (QGNNs) hold significant potential for addressing tasks where classical models face computational bottlenecks.
@@
Early efforts to develop QGNNs have demonstrated promise by extending classical frameworks as Graph Convolutional Networks (GCNs) [15] to the quantum domain via digital quantum circuits [16, 17, 18, 19, 20, 21, 22]. However, despite these advancements, current QGNN frameworks still encounter challenges in scalability and generalization to unseen data. As a result, current quantum approaches struggle to address the broader challenges of real-world graph datasets, where adaptability, scalability, and generalization are crucial for success.
@@
Present Work. To address these challenges, we propose a versatile QGNN framework for inductive node embedding, extending into the quantum realm the classical GraphSAGE approach [23]. This framework samples neighborhood nodes and employs aggregator functions to efficiently combine their feature information. By integrating node features into the learning process, it captures both the topological structure and feature distribution of neighborhoods, enabling generalization to unseen data and scalability. Building on this foundation, our QGNN incorporates a similar sampling mechanism, with quantum circuits now serving as aggregators. The circuit consists of parametrized quantum convolutional and quantum pooling layers, inspired by the Quantum Convolutional Neural Network (QCNN) architecture [24].
To rigorously evaluate the performance of our QGNN framework, we conduct node regression experiments on the QM9 dataset [25, 26]. Molecules are represented as graphs, in which the nodes are corresponding to the atoms and edges to the bonds. Our results demonstrate that the quantum model performs well on the sample data, achieving strong performance metrics. While classical GNNs consistently outperform during training, the quantum model achieves comparable or slightly superior generalization, particularly as molecular complexity increases (i.e., molecules with more atoms). This highlights the robustness of our QGNN framework, which maintains consistent generalizability even for molecules with varying numbers of atoms. Based on the results of this investigation on a non-trivial use-case, we want to motivate further exploration of similar (quantum convolutional) architectures that can potentially go beyond classically simulable problems [27].
Furthermore, we numerically investigate the scalability of the QGNN framework with respect to an increasing number of qubits. This is achieved by addressing the barren plateau problem [28, 29, 30, 31, 32], where gradients vanish exponentially, hindering optimization. Leveraging insights from QCNN research [33, 34], we adopt the same conditions to show the absence of barren plateaus in our framework. Using random input node features, we numerically confirm that the quantum circuit used as an aggregator does not exhibit barren plateaus by showing that the variance of the loss function partial derivatives scales sub-exponentially with the number of qubits. Consequently, since our QGNN employs a single QGCN as an aggregator, we confirm the absence of barren plateaus in the framework, motivating its use for further investigations on more complex graph-based data.
In summary, our work makes three key contributions:
@@
    We evaluate the performance of our proposed approach on a non-trivial dataset (QM9).
@@
    We numerically investigate the QGNN framework’s scalability by confirming the absence of barren plateaus as the number of qubits increases via the variance of the loss gradients.
@@
The paper is structured as follows: Section 2 introduces the problem dataset and explains the node regression task we aim to investigate in this work. Section 3 reviews prior work on QGNNs, identifying their limitations and proposing how our framework addresses these challenges. Section 4 provides a foundational overview of classical GNNs, with a particular focus on the GraphSAGE approach. Section 5 details our proposed QGNN framework, including the design of the quantum circuit and a discussion of its advantages and potential benefits. Section 6 presents the numerical results for both classical and quantum models benchmarked on the QM9 dataset, comparing their performance. We also numerically confirm the absence of barren plateaus in our QGNN, suggesting that the framework is scalable and capable of handling more complex graph data. Finally, Section 7 discusses the results and provides concluding remarks.
2 Molecular Data (QM9)
@@
The QM9 dataset [25, 26] is a standard benchmark in computational chemistry, containing a set of small organic molecules with up to 9 heavy atoms with some being carbon, oxygen, nitrogen, and fluorine. It provides geometric, energetic, electronic, and thermodynamic properties calculated using density functional theory (DFT), including atomization energies, dipole moments, and HOMO-LUMO gaps. Molecular properties such as the LUMO (Lowest Unoccupied Molecular Orbital) energy, which represents the energy level of the lowest unoccupied orbital, and the HOMO (Highest Occupied Molecular Orbital) energy, are key indicators of a molecule’s reactivity and stability. These properties are inherently molecular, meaning they describe the entire molecule rather than individual atoms.
Specifically, we focus on a node regression task, i.e., using atomic properties such as node features to predict molecular properties. In our case, we employ 11 atomic features including amongst them the atomic number, hybridization state, and partial charges to predict the LUMO energy of the molecule. In this way, node regression leverages local atomic information to infer global molecular behavior.
In this investigation, each molecule is represented as a graph, where nodes encode atomic properties, and edges only indicate connectivity between atoms. For each molecule in the QM9 dataset, we extract atom features and the adjacency matrix. However, only the atom features are used as classical input to the model, while the adjacency matrix specifies the set of atom features to be used.
3 Related Work
@@
Quantum Graph Neural Networks have emerged as a promising approach to leverage quantum computing for graph-based learning tasks. Verdon et al. [35] introduced one of the first analog QGNN frameworks, leveraging Hamiltonian evolution for tasks such as learning quantum dynamics and graph isomorphism classification using variational Ansätze. Building on this, Henry et al. [36] extended the concept by applying quantum kernel methods to classification tasks.
@@
In contrast to analog approaches, recent research has shifted toward digital implementations using parameterized quantum circuits (PQC) for solving graph-related problems For instance, Beer et al. [37] developed quantum neural networks (QNNs) for graph-structured quantum data, while Skolik et al. [38] introduced QAOA-type circuits designed for weighted graphs. Additionally, hybrid quantum-classical approaches have emerged as a promising solution to address qubit limitations. Ai et al. [16] proposed a framework that decomposes large graphs into smaller subgraphs, and Zhou et al. [22] combined classical clustering with PQCs. Expanding on hybrid approaches, Tuysuz et al. [39] demonstrated a quantum-classical GNN applied to the reconstruction of the particle track, further showcasing the potential of hybrid methods.
Recent advances have been made in fully quantum graph neural network (QGNN) frameworks, but significant challenges remain. Zheng et al. [20] pioneered the development of the first Quantum Graph Convolutional Network (QGCN) circuits. In their subsequent work [21], the authors proposed a QGCN with convolutional and pooling layers, providing theoretical analysis of the model. They address the barren plateau issue during training and demonstrate promising performance on datasets for graph classification tasks. However, their approach processes the full graph in a single quantum circuit, requiring qubits to encode both the adjacency matrix, nodes and edges. This can potentially lead to increased circuit sizes even for simple graphs. For larger and more complex graphs, the circuit size increases significantly, posing substantial scalability challenges. Similar limitations are observed in Liao et al. [18], whose approach faces scalability issues due to extensive circuitry and lacks numerical benchmarking. Hu et al. [17] proposed a QGCN for semi-supervised node classification tasks, achieving promising results. However, their circuit design, based on edge representations, leads to rapidly increasing circuit depth with the number of edges, creating scalability challenges for dense graphs and potential bottlenecks during training for graphs with specific symmetries. The previous works rely on amplitude encoding to map classical feature vectors to quantum states. Ryu et al. [19] addressed node regression for molecular graphs using equivariantly diagonalizable unitary quantum graph circuits (EDU-QGCs) [40]. However, their atom-to-qubit mapping becomes computationally resource-intensive for larger molecules. Finally, Mernyei et al. [40] explored equivariant quantum graph circuits using small-scale synthetic datasets, suggesting potential scalability challenges when applied to more complex graph structures.
@@
Our QGNN framework adapts GraphSAGE [23] for quantum computation, providing an efficient architecture for inductive learning on diverse graphs. By constructing the graph’s computational representation instead of directly encoding its adjacency matrix into the quantum circuit, we preserve topological features without introducing unnecessary quantum circuit overhead. The quantum circuit integrated into this framework is designed to incorporate flexible convolutional and pooling layers, enabling users to design various Ansätze. Importantly, the number of qubits depends solely on the node feature dimension, keeping the circuit architecture invariant to the number of sampled nodes and their edges. Combined with that, we employ a quantum feature map, instead of amplitude encoding, which offers rich expressivity for reasonably shallow circuit depths[41]. This framework allows classical implementations for heterogeneous graphs, as in Refs. [7, 8, 42, 43], to be extended to the quantum domain. By overcoming the limitations mentioned above, our framework provides a solid foundation for future advancements in quantum graph machine learning.
4 Classical GNN: GraphSAGE
@@
Graph Neural Networks handle graph-structured data, where nodes represent entities and edges capture relationships between them. Unlike traditional neural networks that operate on grid-like data structures, GNNs manage flexible, interconnected structures, making them suitable for diverse applications, including social networks, molecular modeling, and knowledge graphs.
@@
GraphSAGE is a general inductive framework that leverages node feature information to efficiently generate node embeddings. At its core are the sampling and aggregation mechanisms. Using the input graph’s adjacency matrix, the computation graph representation is constructed, preserving local graph structures. Fig. 1 provides a schematic representation of the method.
@@
For a given depth or "hop" of the computational graph, the framework samples a fixed number of neighbors. Embeddings are dynamically generated via a modular aggregation function, producing local neighborhood embeddings for target nodes. Additionally, concatenation can be performed to incorporate the target node’s features with the aggregated neighborhood information. By stacking multiple hops, the target node embedding incorporates information from increasingly larger neighborhoods, capturing multi-hop structural information. This captures neighborhood topology, enabling scalability to large graphs and generalization to unseen data.
(a) Message passing and aggregation mechanisms
(b) Computational graph scheme
Figure 1: Illustration of the GraphSAGE framework. (a) Aggregations (agg1 and agg2) and message passing to the target node $v$ with $k=2$ hops. (b) Transformation of an input graph (left) into a computation graph (right), where nodes represent feature embeddings and edges indicate information flow. Gray boxes denote Multi-Layer Perceptrons (MLPs), which aggregate neighboring embeddings through modular transformations, enabling accurate node representations.
@@
We now delve into the mathematical fundamentals of GraphSAGE, with a focus on the embedding generation and its propagation after the sampling and aggregation steps
4.1 Inner Workings of GraphSAGE

### v2

@@ Abstract
Quantum Graph Neural Networks (QGNNs) offer a promising approach to combining quantum computing with graph-structured data processing. While classical Graph Neural Networks (GNNs) are scalable and robust, existing QGNNs often lack flexibility due to graph-specific quantum circuit designs, limiting their applicability to diverse real-world problems. To address this, we propose a versatile QGNN framework inspired by GraphSAGE, using quantum models as aggregators. We integrate inductive representation learning techniques with parameterized quantum convolutional and pooling layers, bridging classical and quantum paradigms. The convolutional layer is flexible, allowing tailored designs for specific tasks. Benchmarked on a node regression task with the QM9 dataset, our framework, using a single minimal circuit for all aggregation steps, handles molecules with varying numbers of atoms without changing qubits or circuit architecture. While classical GNNs achieve higher training performance, our quantum approach remains competitive and often shows stronger generalization as molecular complexity increases. We also observe faster learning in early training epochs. To mitigate trainability limitations of a single-circuit setup, we extend the framework with multiple quantum aggregators on QM9. Assigning distinct circuits to each hop substantially improves training performance across all cases. Additionally, we numerically demonstrate the absence of barren plateaus as qubit numbers increase, suggesting that the proposed model can scale to larger, more complex graph-based problems.
I Introduction
@@
Graphs provide a versatile framework for modeling complex relationships, with applications ranging from molecular structures to social networks. Unlike traditional machine learning models such as convolutional neural networks (CNNs) and feedforward neural networks (NNs), graph neural networks (GNNs) are specifically designed to exploit the relational structure inherent in graph data. This unique capability enables GNNs to excel in tasks such as node classification, link prediction, and graph regression, where understanding the relationships between entities is as critical as understanding the entities themselves. As a result, GNNs have found widespread use across diverse domains, including social network analysis [borisyuk_2024_lignn, fan_2019_graph], recommendation systems [de_2024_personalized, jain_2019_food], drug discovery [stokes_2020_deep], traffic prediction [derrow_2021_eta], and polypharmacy side effects [kondor_modeling_polypharmacy_2023, zitnik_2018_modeling].
@@
While classical GNNs have achieved remarkable success, the emergence of quantum computing introduces transformative potential for machine learning. Quantum Machine Learning (QML) harnesses principles of quantum mechanics, such as entanglement and superposition to achieve significant speedups over classical computing for specific tasks [arute2019quantum, google2025, cerezo_2022_challenges, coles_2021_seeking, perdomo_2018, ciliberto_2018, liu_2021]. In the context of graph-based problems, Quantum Graph Neural Networks (QGNNs) hold significant potential for addressing graph tasks where classical models may face computational bottlenecks.
@@
Early efforts to develop QGNNs have demonstrated promise by extending classical frameworks as Graph Convolutional Networks (GCNs) [welling_GCN_2017] to the quantum domain via digital quantum circuits [hu_qgnn_2022, yidong_qgnn_2024, ryu_qgnn_2023, zheng_qgnn_2021, zheng_qgnn_2024]. However, despite these advances, the existing QGNN frameworks still face significant limitations in scalability and in their ability to generalize to unseen data. This is largely because their quantum circuits are tailored to fixed graph structures, preventing them from adapting to the diversity of real-world graph datasets. As a result, current quantum approaches struggle to meet the demands of practical applications, where adaptability, scalability, and strong generalization are essential.
@@
Present Work. To address these challenges, we propose a versatile QGNN framework for inductive node embedding, extending into the quantum realm the classical GraphSAGE approach [leskovec_inductive_representation_2018]. This framework samples uniformly neighborhood nodes and uses quantum circuits as aggregators to efficiently combine their features in an inductive manner, enabling both scalability and generalization to unseen data. The circuit is composed of parameterized quantum convolutional and pooling layers, inspired by the Quantum Convolutional Neural Network (QCNN) architecture [lukin_qcnn_2019].
To rigorously evaluate the performance of our QGNN framework, we conduct node regression experiments on the QM9 dataset [pyg_code, moleculeNet_2017]. Molecules are represented as graphs, in which the nodes are corresponding to the atoms and edges to the bonds. Our results demonstrate that the quantum model performs well on the sample data, achieving strong performance metrics. However, when using a single quantum circuit reused across all aggregation steps, classical GNNs consistently outperform the quantum model during training, whereas the quantum model exhibits stronger generalization, particularly as molecular size increases (i.e., for molecules with a larger number of atoms). This highlights the robustness of our QGNN framework in maintaining consistent generalization across molecules of varying sizes, despite relying on a minimal model. The aforementioned trainability limitations of the single-circuit setup are addressed by introducing multiple independent quantum aggregators, leading to higher training performance of the quantum approach. Furthermore, we investigate the scalability of the QGNN framework with respect to an increasing number of qubits. This is achieved by addressing the barren plateau problem [anschuetz_2022_beyond, ragone_2023_lie, larocca_2024_review, arrasmith_2021_equivalence], where gradients vanish exponentially, hindering optimization. Leveraging previous results [ragone_2023_lie] and using random input node features, we numerically show that the proposed QGNNs do not exhibit barren plateaus since the variance of the loss function partial derivatives scales sub-exponentially with the number of qubits. This suggests that our framework is scalable and capable of handling complex graph data, motivating its use for further investigations on more complex graph-based data.
In summary, our work makes four key contributions:
@@
    We evaluate our approach on QM9 using a single minimal QGCN circuit, reused at every aggregation step, to show that it can handle graphs with varying numbers of nodes without changing the number of input qubits.
@@
    We improve the trainability of QGNNs by extending the framework incorporate multiple quantum aggregators.
  4. 4.
@@
    We numerically investigate the QGNN framework’s scalability by confirming, via the variance of the loss gradients, the absence of barren plateaus as the number of qubits increases.
@@
The paper is structured as follows: Section II introduces the problem dataset and explains the node regression task we aim to investigate in this work. Section III reviews prior work on QGNNs, identifying their limitations and proposing how our framework addresses these challenges. Section IV provides a foundational overview of classical GNNs, with a particular focus on the GraphSAGE approach. Section V details our proposed QGNN framework, including the design of the quantum circuit and a discussion of its advantages and potential benefits. Section VI presents the numerical results for both classical and quantum models benchmarked on the QM9 dataset, comparing their performance. We also numerically confirm the absence of barren plateaus in our QGNN. Finally, Section VII discusses the results and provides concluding remarks.
II Molecular Data (QM9)
@@
The QM9 dataset [pyg_code, moleculeNet_2017] is a standard benchmark in computational chemistry, comprising small organic molecules built from combinations of nine heavy atoms, including carbon, oxygen, nitrogen, and fluorine. It provides geometric, energetic, electronic, and thermodynamic properties calculated using density functional theory (DFT), including atomization energies, dipole moments, and HOMO-LUMO gaps (Highest-Lowest Unoccupied Molecular Orbital). In particular, the LUMO and HOMO energies are key indicators of a molecule’s reactivity and stability. These properties are inherently molecular, describing the molecule as a whole rather than individual atoms.
Specifically, we focus on a node regression task, i.e., using atomic properties such as node features to predict molecular properties. The dataset provides 11 atomic features, but we use only 7 of them: atomic number, chirality, degree, formal charge, radical electrons, hybridization, and scaled mass. The remaining four (aromaticity, hydrogen count, ring membership, and valence) are excluded since they are consistently zero in the selected samples used in this work. In this way, node regression leverages local atomic information to infer global molecular behavior. Essentially, in this investigation, each molecule is represented as a graph, where nodes represents atoms, and edges only indicate connectivity between atoms. For each molecule in the QM9 dataset, we extract atom features and the adjacency matrix. However, only the atom features are used as classical input to the model, while the adjacency matrix specifies the set of atom features to be used as input.
III Related Work
@@
Quantum Graph Neural Networks have emerged as a promising approach to leverage quantum computing for graph-based learning tasks. Verdon et al. [verdon_2019_quantumgraph] introduced one of the first analog QGNN frameworks, leveraging Hamiltonian evolution through global unitary dynamics for tasks such as Hamiltonian learning, entanglement generation, and graph isomorphism classification. Building on this, Henry et al. [thabet_2024] extended the concept by applying quantum kernel methods to classification tasks.
@@
In contrast to analog approaches, a few works have shifted toward digital implementations using variational circuits. For instance, Beer et al. [beer_2021_quantum] developed quantum neural networks for graph-structured quantum data, while Skolik et al. [skolik_2023_equivariant] introduced QAOA-type circuits designed for weighted graphs. Hybrid quantum–classical approaches have also emerged as possible alternatives to constraints posed by quantum circuits. Ai et al. [xing_qgnn_2024] proposed a framework that decomposes large graphs into smaller subgraphs, and Zhou et al. [kaixiong_qgnn_2022] combined classical clustering with variational circuits. Expanding on hybrid approaches, Tuysuz et al. [tuysuz_2021_hybrid] demonstrated a quantum-classical GNN applied to the reconstruction of the particle track, further developing hybrid methods.
Recent progress has been made in fully digital QGNNs, yet major challenges remain. Zheng et al. [zheng_qgnn_2021, zheng_qgnn_2024] introduced QGCN architectures with convolution and pooling layers, addressing barren plateaus and reporting strong performance. However, their design encodes the full graph—adjacency, nodes, and edges—within a single quantum circuit, causing circuit width and depth to scale rapidly with graph size. Liao et al. [yidong_qgnn_2024] face similar overheads due to large circuit constructions and lack numerical benchmarks. Hu et al. [hu_qgnn_2022] design an edge-based QGCN for semi-supervised node classification, where circuit depth can increase rapidly with the number of edges, potentially resulting in large circuits and training bottlenecks for graphs with particular symmetries. All these approaches rely on amplitude encoding. Ryu et al. [ryu_qgnn_2023] employ equivariantly diagonalizable unitary quantum graph circuits (EDU-QGCs) for node regression in molecules, but the atom-to-qubit mapping becomes costly for larger molecules. Finally, Mernyei et al. [mernyei_2022_equivariant] evaluate equivariant quantum graph circuits only on small synthetic datasets, leaving scalability to more complex graphs uncertain.
@@
Conversely, our QGNN framework adapts GraphSAGE [leskovec_inductive_representation_2018] for quantum computing, providing an efficient architecture for inductive learning on diverse graphs. By constructing the graph’s computational representation instead of directly encoding its adjacency matrix into the quantum circuit, we preserve topological features without introducing unnecessary quantum circuit overhead. The circuit integrated into this framework is designed to incorporate flexible convolutional and pooling layers, enabling users to design various Ansätze. Importantly, the number of qubits depends solely on the node feature dimension, keeping the circuit architecture invariant to the number of sampled nodes and edges. Combined with that, we employ a nonlinear quantum feature map, instead of amplitude encoding, which offers rich expressivity for reasonably shallow circuit depths [kyriienko_2021_solving]. By addressing the limitations outlined above, our framework offers a solid foundation for advancing quantum graph machine learning, including potential extensions of classical heterogeneous-graph implementation to the quantum domain, as in Refs. [kondor_modeling_polypharmacy_2023, zitnik_2018_modeling, leskovec_embedding_2019].
IV Classical GNN: GraphSAGE
@@
GNN handle graph-structured data, where nodes represent entities and edges capture relationships between them. Unlike traditional neural networks that operate on grid-like data structures, GNNs manage flexible, interconnected structures, making them suitable for diverse applications, including social networks, molecular modeling, and knowledge graphs.
@@
GraphSAGE is a general inductive framework that leverages node feature information to efficiently generate node embeddings. At its core are the sampling and aggregation mechanisms. Using the input graph’s adjacency matrix, the computation graph representation is constructed, preserving local graph structures. Fig. 1 provides a schematic representation of the method, where in particular the message generation and passing are illustrated in Fig. 1(a).
@@
For a given depth or ”hop” of the computational graph, the framework samples a fixed number of neighbors. Embeddings are dynamically generated via a modular aggregation function, producing local neighborhood embeddings for target nodes. Additionally, concatenation can be performed to incorporate the target node’s features with the aggregated neighborhood information. By stacking multiple hops, the target node embedding incorporates information from increasingly larger neighborhoods, capturing multi-hop structural information. This captures neighborhood topology, enabling scalability to large graphs and generalization to unseen data.
@@
We now delve into the mathematical fundamentals of GraphSAGE, with a focus on the embedding generation and its message after the sampling and aggregation steps
(a) Message passing and aggregation mechanisms.
(b) Computational graph scheme.
Figure 1: Illustration of the GraphSAGE framework. (a) Aggregation steps: agg1 and agg2, generate the message which is then passed to the target node $v$ for $k=2$ hops. (b) Transformation of an input graph (left) into a computation graph (right), where nodes represent feature embeddings and edges indicate information flow. Gray boxes denote Multi-Layer Perceptrons (MLPs), which aggregate neighboring embeddings through modular transformations, enabling accurate node representations.
IV.1 Inner Workings of GraphSAGE
In the GraphSAGE framework, the GNN generates node embeddings through two main stages: neighborhood sampling and feature aggregation, where the aggregation is performed by an MLP-based aggregator. In particular, the MLP are defined by an inductive variant of the GCN. For brevity, we refer to this aggregator simply as GCN throughout the text.